OpenStudy (anonymous):

What is the value of x in the equation 4x + 8y = 40, when y = 0.8? 4.6 8.4 10 1

OpenStudy (photon336):

you could do this two ways you can solve for x and then plug in y, do you know how to do this?

OpenStudy (photon336):

subtract 8y from both sides to get this \[40-8y = 4x\]

OpenStudy (photon336):

divide both sides by 4 \[(40-8y)/(4) = x \] this gives you 10-2y = x

OpenStudy (photon336):

plug in y = 0.8

OpenStudy (photon336):

10-2(8/10) = 8.4 does this make sense? @apalmer004 ?

OpenStudy (anonymous):

Yes, thank you.

OpenStudy (photon336):

excellent, no problem